# FauForensics: Boosting Audio-Visual Deepfake Detection with Facial Action Units

### v1

@@ IV-E Ablation Study
In this section, we conduct two groups of ablation experiments to analyze the effect of the core module and the feature extractor.
Core Modules. To systematically evaluate the contributions of individual modules, we conduct an ablation study comparing three variants with our proposed method, as shown in  IV. Specifically, our framework contains three core modules: 1) FAU-enhanced Feature Learning (FFL) module, 2) Implicit Feature Alignment (IFA) module, and 3) Temporal Attentional Pooler (TAP). We conduct ablation experiments by training in LAV-DF and testing in FakeAV and LAV-DF, which can be used for performance comparisons in both within-database and cross-database evaluations. Quantitative results reveal that each module can contribute to better detection performance.
@@
In Table IV, we can see that our method and its variants all obtain promising results in within-database evaluation. It should be noted that each module gradually improves the performance of cross-dataset evaluation. Specifically, Variant-B integrated with the FFL module improves about 2.04% AUC (i.e., 79.33 v.s. 77.29) compared with Variant-A in cross-database evaluation. This can illustrate that the FFL module can enhance the discriminative feature learning and promote the performance improvement of audio-visual deepfake detection. In addition, benefit from the IFA module, Variant-C perform implicit feature alignment for better inter-modal learning and obtain 1.64% AUC improvement over Variant-B. Finally, TAP centers on fine-grained frame-wise audiovisual similarities, further boosting audio-visual deepfake detection.
Feature Extractor. To further evaluate the effectiveness of each feature extractor in our proposed method, we compare our method with three variants, as shown in Table IV. The quantitative results demonstrate the importance of each feature extractor. Notably, the variant without the FAU encoder suffers from significant performance degradation in cross-dataset evaluation, quantitatively confirming the essential role of facial action units in our proposed framework.
@@
Figure 5: The t-SNE visualization of features derived from our proposed method and ours w/o FAUs on the test set of within-database and cross-database. Red, blue, green, and violet indicate $RARV$, $RAFV$, $FARV$, and $FAFV$, respectively.
@@ IV-F Visualization Analysis
To further illustrate the effectiveness of FAUs, we utilize t-distributed stochastic neighbor embedding (t-SNE) [43] to viusalize features derived from our proposed method and ours w/o FAUs on the test set of within-database and cross-database (training on FakeAV), as depicted in Fig. 5. It can be observed that both methods obtain promising in within-database scenarios, but features of ours are more separable than the that of ours w/o FAUs in cross-database scenarios. The circled part of Fig. 5(c) shows the weakened discriminability due to the lack of FAUs. Ours without FAUs can not distinguish $RARV$ (red part in Fig. 5) and $RAFV$ (blue part) well. However, ours benefits from FAUs can force the model to focus on richer facial motion cues and show discriminative capacity for audio-visual deepfake detection. Therefore, the above observation demonstrates that FAUs can boost feature learning of detectors.
@@ V Conclusion
In this work, we propose a novel FAU-guided framework with frame-level integration to simultaneously capture various forgeries (i.e., audio, visual and multi-modal forgeries). More precisely, we firstly introduce FAUs to learn facial geometric information for capturing FAU-related forgeries. Moreover, frame-level integration is designed to capture intra- and inter-modal temporal correlation from audio-visual signals. Extensive experiments on two widely used audio-visual deepfake databases illustrate the superiority of our proposed method to prior SOTA methods.
In future work, we will delve into finer-grained audio features (e.g., phonemes) and leverage LLMs to detect temporal inconsistencies between phonemes and facial action units (FAUs), thereby enhancing generalization capabilities and interpretability of forgery detectors.

### v2

@@ IV-E Ablation Study
In this section, we conduct two groups of ablation experiments to analyze the effect of the core modules and the feature extractors.
Core Modules. To systematically evaluate the contribution of each core module in our framework, we conduct an ablation study by comparing three variants with the full proposed method: Variant-A, Variant-B, and Variant-C, as summarized in Table IV. Specifically, our framework consists of three key modules: (1) FAU-enhanced Feature Learning (FFL) module, (2) Implicit Feature Alignment (IFA) module, and (3) Temporal Attentional Pooler (TAP). The ablation experiments are performed by training on the LAV-DF dataset and testing on both FakeAV and LAV-DF, enabling evaluation under both within-database and cross-database scenarios. The quantitative results clearly demonstrate that each module contributes to overall performance improvement, validating the effectiveness of our designed module.
@@
In Table IV, we can see that both our method and its variants obtain promising results in within-database evaluations. It should be noted that each module gradually improves the performance of cross-dataset evaluation. Specifically, Variant-B, which incorporates the FAU-enhanced Feature Learning (FFL) module, achieves about 2.04% AUC (i.e., 79.33 vs. 77.29) improvement over Variant-A in cross-database evaluation. This can illustrate that the FFL module can enhance the discriminative feature learning and promote the performance improvement of audio-visual deepfake detection. In addition, benefit from the IFA module, Variant-C performs implicit feature alignment for better inter-modal learning and obtains 1.64% AUC improvement over Variant-B. Finally, TAP centers on fine-grained frame-wise audiovisual correlations, further boosting detection performance and contributing to the overall robustness of our approach.
Feature Extractors. Our proposed architecture contains three feature extractors: a FAU encoder, a Video encoder, and an Audio encoder. To evaluate the effectiveness of each feature extractor in our architecture, we compare our full model with three variants, as presented in Table V. The quantitative results clearly demonstrate the contribution of each feature extractor to the overall performance. In particular, the variant without the FAU encoder suffers from significant performance degradation in cross-dataset evaluations, quantitatively confirming the essential role of facial action units in enhancing the generalization capability of our architecture.
@@
Figure 7: The t-SNE visualization of features derived from our proposed method and ours w/o FAUs on the test sets of both within-database and cross-database evaluations. Red, blue, green, and violet indicate $RARV$, $RAFV$, $FARV$, and $FAFV$, respectively.
[…]
@@ IV-F Visualization Analysis
To further illustrate the effectiveness of FAUs, we utilize t-distributed stochastic neighbor embedding (t-SNE) [60] to visualize features derived from our proposed method and ours w/o FAUs on the test set of within-database and cross-database (training on FakeAV), as depicted in Fig. 7. It can be observed that both methods obtain promising results in within-database scenarios, but the features of ours are more separable than those of ours w/o FAUs in cross-database scenarios. The circled part of Fig. 7(c) shows the weakened discriminability due to the lack of FAUs. Ours without FAUs can not distinguish $RARV$ (red part in Fig. 7) and $RAFV$ (blue part) well. However, our method benefits from the incorporation of FAUs, which can force the model to focus on more fine-grained facial motion cues, thereby enhancing its discriminative capability for audio-visual deepfake detection. Therefore, the above observation demonstrates that FAUs can boost feature learning of detectors.
To better understand the role of FAU in the audio-visual forgery detection task, we visualize the phoneme-FAU pairings from four groups of audio-visual signals in Fig. 8. Facial action units describe nearly all possible facial actions. For example, AU25 (lip part) can depict the mouth motion, which is closely related to pronunciation. Typically, the pronunciation of certain phonemes is closely linked to the shape of the mouth. For instance, the pronunciation of the phoneme $/\rm p/$ requires the mouth to be tightly closed first. Therefore, we focus on the phoneme $/\rm p/$ and AU25 to show frame-level audio-visual forgery cues. In Fig. 8, we observe that real audio-visual signals display the correct pronunciation of $/\rm p/$ in the word “pulled”, which is to close the mouth first and then open it. However, the forged visual signals do not exhibit appropriate mouth movements corresponding to the phoneme at that moment for the RAFV pairing. Similarly, FARV did not follow this pronunciation mechanism. Both sets of visual signals indicate that the mouth remains slightly open, which does not match the pronunciation of $/\rm p/$. Finally, although FAFV improves lip-syncing via wavtolip, subtle traces imperceptible to the human eye can still be found in the frame-by-frame analysis of phoneme-FAU. For example, FAFV also exhibits a close-to-open mouth motion, with the mouth closure delayed by one audio frame.
@@ V Conclusion
In this work, we propose a novel FAU-guided framework with frame-level integration to simultaneously capture various forgeries (i.e., audio, visual, and multi-modal forgeries). More precisely, we first introduce FAUs to learn facial geometric information for capturing FAU-related forgeries. Moreover, a frame-level integration module is designed to capture intra- and inter-modal temporal correlations from audio-visual signals. Extensive experiments on two widely used audio-visual deepfake databases illustrate the superiority of our proposed method to prior SOTA methods.
Limitations. Despite the promising performance, our method has some limitations. For example, our detection framework relies on the availability of both audio and visual streams. This requirement also restricts the method’s applicability to scenarios involving incomplete or single-modal media, such as GIF and muted videos. Moreover, our approach requires that the audio is clearly aligned with the on-screen speaker, as audio originating from off-screen individuals or background music can disrupt the modeling of audio-visual correspondence and negatively impact detection accuracy.
Future work. To address the limitations above, future work will explore three main directions. Firstly, we plan to leverage large language models (LLMs) to develop a more flexible detection framework capable of handling incomplete or single-modal content. Secondly, to mitigate the reliance on aligned audio-visual inputs, we aim to incorporate speaker localization and audio source separation techniques. Finally, we will attempt to delve into finer-grained audio features (e.g., phonemes) and utilize LLMs to detect temporal inconsistencies between phonemes and facial action units (FAUs), thereby enhancing generalization capabilities and interpretability of forgery detectors.